OpenStudy (jt950):

If gradient of tangent=sin x -(3/4), sketch the graph of the curve

OpenStudy (jt950):

(0,5pi)

OpenStudy (dumbcow):

is this saying \[\frac{dy}{dx} = \sin x - \frac{3}{4}\] ?

OpenStudy (jt950):

ya

OpenStudy (dumbcow):

oh ok then to get curve of f(x), take integral use given point to get constant of integration

OpenStudy (dumbcow):

int of sin = -cos

OpenStudy (jt950):

can draw a graph for me? thx

OpenStudy (dumbcow):

a graph of what? use wolfram http://www.wolframalpha.com/input/?i=plot+sin%28x%29+-.75

OpenStudy (jt950):

thank you